# From Estimands to Robust Inference of Treatment Effects in Platform Trials

### v1

@@ 8 Summary and Recommendations
Platform trials offer the potential to enhance statistical efficiency by facilitating the sharing of data between different treatment evaluations. However, realizing this potential requires careful consideration of several challenges, such as defining meaningful estimands that reflect eligibility and concurrency, and developing robust estimation techniques that accommodate the varying probabilities of treatment assignments over time and differing patient characteristics.
In this article, we explicitly present how to construct a clinically meaningful estimand with clear specificity on the population of concurrently eligible patients. To estimate and infer this estimand with minimal statistical assumptions, we develop methods using weighting or post-stratification to account for varying treatment assignment probabilities. We also consider model-assisted approaches for covariate adjustment. We provide explicit formulas for robust variance estimation, which have been implemented in our R package, RobinCID. Through theoretical analysis and simulation studies assessing the relative efficiency of these methods, we find that the stabilized IPW (SIPW) and post-stratification (PS) significantly enhance the efficiency of the IPW estimator. Additionally, model-assisted approaches for covariate adjustment (SAIPW or APS) generally improve efficiency while maintaining robustness against model misspecification. A detailed relative efficiency comparison is in Figure 3. Note that although SIPW is less efficient than PS asymptotically, it is still useful when there are sparse strata because SIPW may have better finite-sample performance and there may be situations where the PS estimator or its variance estimator cannot be computed.

### v4

@@ 7 Summary and Recommendations
Master protocol trials have the potential to enhance statistical efficiency by facilitating the sharing of data across treatment evaluations. However, realizing this potential requires careful consideration of two fundamental challenges: defining meaningful estimands to reflect eligibility and concurrency, and developing robust estimation methods to accommodate the treatment assignment probabilities varying with time and individual characteristics.
In this article, we present a clear framework for constructing a clinically meaningful estimand with precise specification of the population. Our main discussion centers on comparing two treatments within the ECE population, as it preserves the integrity of randomization and avoids arbitrary dependence on randomization probabilities. This framework is adaptable to various contexts, such as comparisons across multiple treatments and evaluation of treatment effects within specific subgroups or a re-weighted population. Achieving clarity on the estimand is crucial for addressing other key issues in future statistical research, including interim analyses and the use of nonconcurrent controls. To estimate and infer this estimand with minimal statistical assumptions, we develop methods using weighting or post-stratification to account for varying treatment assignment probabilities. We also consider model-assisted approaches for covariate adjustment. Theoretical and empirical results are given and a detailed relative efficiency comparison is in Figure 5. We provide explicit formulas for robust variance estimation, which have been implemented in our R package, RobinCID. Lastly, our results are applicable not only to master protocol trials but also more broadly to trials with constrained or non-uniform treatment assignment.